# On How Iterative Magnitude Pruning Discovers Local Receptive Fields in Fully Connected Neural Networks

### v1

@@ Abstract
Since its use in the Lottery Ticket Hypothesis, iterative magnitude pruning (IMP) has become a popular method for extracting sparse subnetworks that can be trained to high performance. Despite this, the underlying nature of IMP’s general success remains unclear. One possibility is that IMP is especially capable of extracting and maintaining strong inductive biases. In support of this, recent work has shown that applying IMP to fully connected neural networks (FCNs) leads to the emergence of local receptive fields (RFs), an architectural feature present in mammalian visual cortex and convolutional neural networks. The question of how IMP is able to do this remains unanswered. Inspired by results showing that training FCNs on synthetic images with highly non-Gaussian statistics (e.g., sharp edges) is sufficient to drive the formation of local RFs, we hypothesize that IMP iteratively maximizes the non-Gaussian statistics present in the representations of FCNs, creating a feedback loop that enhances localization. We develop a new method for measuring the effect of individual weights on the statistics of the FCN representations (“cavity method”), which allows us to find evidence in support of this hypothesis. Our work, which is the first to study the effect IMP has on the representations of neural networks, sheds parsimonious light one way in which IMP can drive the formation of strong inductive biases.
@@ 1 Introduction
Iterative magnitude pruning (IMP) [1] has emerged as a powerful tool for identifying sparse subnetworks (“winning tickets”) that can be trained to perform as well as the dense model they are extracted from [2, 3]. That IMP, despite its simplicity, is more robust in discovering such winning tickets than other, more complex pruning schemes [4] suggests that its iterative coarse-graining [5] is especially capable of extracting and maintaining strong inductive biases. This perspective is strengthened by observations that winning tickets discovered by IMP: 1) have properties that make them transferable across related tasks [6, 7, 8, 9, 10, 11, 12, 13] and architectures [14]; 2) can outperform dense models on classes with limited data [15]; 3) have less overconfident predictions [16].
@@
The first direct evidence for IMP discovering good inductive biases came from studying the winning tickets extracted by IMP in fully connected neural networks (FCNs) [17]. Pellegrini and Biroli (2022) [17] found that the sparse subnetworks identified by IMP had local receptive field (RF) structure (Fig. 1A), an architectural feature found in visual cortex [18] and convolutional neural networks (CNNs) [19]. Comparing IMP derived winning tickets with the sparse subnetworks found by oneshot pruning (Fig. 1B), Pellegrini and Biroli (2022) [17] argued that the iterative nature of IMP was essential for refining the local RF structure. However, to-date, an understanding of how IMP, a pruning method based purely on the magnitude of the network parameters, is able to “sift out” non-localized weights remains unknown. Resolving this will not only shed light on the effect of IMP on FCNs, but also will provide new insight on the success of IMP more broadly.
@@
Figure 1: IMP discovers more localized RFs than oneshot magnitude pruning of FCNs. (A) Localized RFs are present after applying IMP for 10 rounds of pruning (each round pruning $s=30\%$ of the remaining weights), leading to a subnetwork with $s=97.2\%$. (B) Noisier, less localized RFs are present in the masks found after oneshot pruning FCNs trained on ImageNet32 to $s=97.2\%$ sparsity. Pruned weights are shown in black and remaining weights are colored by which input channel (red, green, blue) they are connected to. The masks shown correspond to the hidden units with the greatest number of weights remaining [17].
@@
Historically, the study of local RFs has focused on specific features of natural images (e.g., sharp edges), and has shown that, with regularization, it is possible for local RFs to emerge in FCNs [20, 21, 22, 23]. Recent work has built upon this, showing that synthetic images with sufficiently strong non-Gaussian statistics can, without any regularization, drive the formation of local RFs in FCNs [24, 25]. Inspired by this, we hypothesize that IMP (Sec. 2.1) is able to discover local RFs (Sec. 2.2) in FCNs by iteratively increasing the non-Gaussian statistics present in its internal representation. This could create a feedback loop, where the amplification of non-Gaussian statistics leads to greater localization, which leads to further increases in non-Gaussian statistics. Because of IMP’s broad success in computer vision, we hypothesize that IMP maximally increases the non-Gaussian statistics, by removing exactly the weights that maximize the “non-Gaussanity” of the internal representations.
@@
Hypothesis (H): IMP discovers local RFs in FCNs by maximally increasing the non-Gaussian statistics represented by the network.
@@
While this is a challenging hypothesis to prove, we provide the following evidence in support of $\textbf{H}^{*}$:
@@
  • Training FCNs on “Gaussian clones” [26] that match the original dataset’s first two moments (mean and covariance), but have no higher-order moments, we find that IMP does not discover localized RFs. This demonstrates that non-Gaussian statistics are necessary for IMP to discover local RFs in FCNs (Sec. 3.1).
@@
  • Measuring the non-Gaussian statistics present in the preactivations of the FCN, we find that IMP leads to representations with greater non-Gaussian statistics, as compared to oneshot pruning (Sec. 3.2). This difference increases with each round of pruning, demonstrating the importance of the iterative nature of IMP.
@@
  • Developing a method that enables measuring the effect individual weights have on the statistics of the FCN representations, we find that IMP removes weights precisely when their pruning would increase the non-Gaussanity of the FCN representations (Sec. 3.3). This suggests the optimality of IMP in not just increasing, but maximizing, the non-Gaussian statistics at each round of pruning.
@@
Collectively, our work provides the first in-depth analysis of how IMP affects the internal representations of DNNs, and provides insight on how the statistics of internal representations interact with the iteratively evolving architecture achieved by IMP. Additionally, it highlights the limitations of theories for that have been developed assuming Gaussian representations [27, 28, 29, 30, 31, 32, 33], which remain dominant in the theoretical machine learning community. We hope our work motivates future exploration to reveal IMP’s potential for learning parsimonious structure in DNNs, and believe the tools we make use of (Gaussian data clones and the cavity method) offer new ways of achieving this goal.
@@ 2.1 Iterative magnitude pruning
Given a neural network $f(\theta;X)$, where $\theta$ are its $N$ parameters (e.g., weights, biases) and $X$ is a set of data samples used for training that come from an underlying distribution $\mathcal{X}$, IMP [1, 2] performs the following iterative coarse-graining procedure. First, $f(\theta,X)$ is trained for $T$ iterations, resulting in $f[\theta(T);X]$. Then, a mask $m\in\{0,1\}^{N}$ is computed by assigning $m_{i}=0$ if $|\theta_{i}(T)|<\tau$, for some threshold $\tau\in\mathbb{R}^{+}$. For network parameters with magnitude greater than $\tau$, $m_{i}=1$. Typically, $\tau$ is set such that $s=1-\frac{1}{N}\sum_{i=1}^{N}m_{i}$, for a desired sparsity $s\in(0,1]$.
Having computed the mask, the network parameter values are “rewound” [34, 35] to a previous value, $\theta(t_{\text{rewind}})$, where $t_{\text{rewind}}<<T$. In the original work leveraging IMP to discover winning tickets, it was demonstrated that winning tickets could be found at initialization (i.e., $t_{\text{rewind}}=0$). Subsequent work has found it necessary to set $t_{\text{rewind}}>0$ [3], particularly when considering architectures larger than LeNet [19] and datasets more complex than MNIST. A common approach is to set $t_{\text{rewind}}\approx 0.01T$. The network $f[\theta(t_{\text{rewind}})\odot m;X]$ is then trained for $T-t_{\text{rewind}}$ training iterations, where $\odot$ denotes element-wise multiplication.
This train, prune, and rewind procedure is then repeated, with round $n$ of IMP involving training $f[\theta(t_{\text{rewind}})\odot m(n-1);X]$ and computing the mask $m(n)\in\{0,1\}^{N}$ from the remaining (non-pruned) parameters $\theta(T)\odot m(n-1)$. The masks $m(n)$ are non-trainable parameters, and $\sum_{i=1}^{N}m_{i}(n)<\sum_{i=1}^{N}m_{i}(n-1)$. This is repeated for $N_{\text{IMP}}$ rounds.
@@ 2.2 Local receptive fields
Local RFs were first discovered in mammalian primary visual cortex (V1) [18], where individual V1 neurons respond to specific visual features (e.g., the presence of lines) in a local range of visual space. The computational paradigm of breaking input space into local patches, over which a hierarchical representation could be learned, inspired the development of CNNs [19].
Early work studying how local RFs might emerge in artificial neural networks showed that $L_{1}$ regularization of hidden unit activations was capable of driving localization in FCNs [20, 21, 22, 23]. Thus, while sparsity in activations was known to lead to local RFs, it was not until subsequent effort using a variant of LASSO regression ($\beta$-LASSO) on network parameters that it was appreciated that sparsity in the weights could also lead to localized RFs [47]. Unlike this prior literature, our work focuses on the emergence of local RFs, without explicit regularization [17, 24, 25].
To quantify the locality of the IMP masks, $m(n)$, we use the following metric. Let $X$, the input to an FCN $f(\theta;X)$, be a square image with $N_{p}^{2}\in\mathcal{N}$ pixels ($N_{p}$ in the $x-$dimension and $N_{p}$ in the $y-$dimension). Let the locations of two pixels of $X$ be denoted as $z=(x,y)$ and $z^{\prime}=(x^{\prime},y^{\prime})$, where $z,z^{\prime}\in[1,...,N_{p}]\times[1,...,N_{p}]$. Let the relative position of $z$ and $z^{\prime}$ be denoted as $d_{zz^{\prime}}=z-z^{\prime}$. We define the number of pixels in $X$, with relative position $d$, that are connected to unit $i$ in the first hidden layer of an FCN, at IMP round $n$, by
@@
where $\delta(d_{zz^{\prime}}-d)$ is a delta function that is equal to $1$ when $d_{zz^{\prime}}=d$, and 00 otherwise. $S_{i}(d,n)$ acts as a correlation function on the mask associated with unit $i$, $m_{i}(n)$. Pelligrini and Biroli (2022) [17] previously used $S(d,n)$ to visualize the locality of the IMP masks. We further use it to quantify the localization by computing the standard deviations, $\sigma_{x}$ and $\sigma_{y}$, of a two-dimensional Gaussian fit to $S_{i}(d,n)$. These standard deviations act as a measure of correlation length, with the smaller $\sigma_{x}$ and $\sigma_{y}$ are, the more localized the RFs of the IMP masks are. We report $\sigma_{x}$ as the RF width (similar results are found with $\sigma_{y}$ ).
@@ 3.1 Non-Gaussian statistics are necessary for IMP to discover local RFs in FCNs
Local information is contained in the non-Gaussian statistics. Natural images have a number of defining properties. One such property, sharp luminosity changes – which feature contours and edges – has been recognized as a hallmark of natural images [20, 21, 22, 23]. From a statistical perspective, these sudden changes translate into large higher-order cumulants (i.e., beyond mean and covariance). Thus, natural images can be said to contain non-Gaussian statistics, as a generative model which samples only from a Gaussian distribution, where all cumulants after the second are zero, cannot reproduce these features.
Previous seminal work on learning local RFs has emphasized the importance of these non-Gaussian statistics [20, 21, 22, 23], and recent work has shown that synthetic images, with strongly non-Gaussian statistics, can be sufficient to drive local RF formation in FCNs [24]. A full dynamical theory of this phenomenon has been developed in a simplified FCN setting [25]. To illustrate why non-Gaussian statistics play a crucial role in the learning of local RFs, we perform independent component analysis (ICA) [48] and principle component analysis (PCA) on images from the ImageNet32 dataset [49], a downsampled version of the original ImageNet dataset (see Appendix A for more details). ICA looks for linear projections, $\lambda=W\cdot X$, which maximize the “non-Gaussianity” of the features $\lambda$. Running this algorithm on ImageNet32 leads to components, $W$, that are localized and resemble Gabor filters (Fig. 2A), a property found in the response tuning of V1 neurons [18]. In contrast, running PCA, which only acts on the covariance of the inputs (and thus, only has access to the Gaussian features of the data), yields $W$ with oscillating components (Fig. 2B). These oscillations provide a spatially distributed and thus, non-local, representation of the images. Similar results have been documented on the full ImageNet dataset [^1] See https://twitter.com/pierreablin/status/1435743782285283330.. Thus, local information is carried primarily by the non-Gaussian statistics of natural images.
@@
Figure 2: Non-Gaussian statistics contain local information in ImageNet32. (A) By maximizing the non-Gaussanity of a lower dimensional representation of the 50,000 validation images from ImageNet32, ICA extracts features that are localized. (B) In contrast, considering only the covariance of the validation images from ImageNet32 leds PCA to extract features that are non-local.
@@
IMP does not discover local RFs when applied to FCNs trained on data with only Gaussian statistics. Given the results from Fig. 2, we hypothesize that non-Gaussian statistics are necessary for IMP to find local RFs in FCNs. To test this, we generate a Gaussian clone of ImageNet32 (ImageNet32 sampled from a Gaussian process – “ImageNet32-GP”) [50]. This dataset matches the mean and covariance of ImageNet32, but is explicitly constructed to not contain any higher-order cumulants (Fig. 3A – see Appendix B.1, for details). Unlike when IMP is applied to FCNs trained on ImageNet32, IMP fails to find local RFs in the ImageNet32-GP setting. The masks contain no obvious structure (Fig. 3B) and our metric for localization (Sec. 2.2) shows that the masks found on ImageNet32-GP have a significantly larger correlation length than the masks found on ImageNet32.
@@
These results demonstrate that non-Gaussian statistics, in the form of higher-order cumulants, are necessary for local RFs to emerge via IMP.

### v3

# How Iterative Magnitude Pruning Discovers Local Receptive Fields in Fully Connected Neural Networks

@@ Abstract
Since its use in the Lottery Ticket Hypothesis, iterative magnitude pruning (IMP) has become a popular method for extracting sparse subnetworks that can be trained to high performance. Despite its success, the mechanism that drives the success of IMP remains unclear. One possibility is that IMP is capable of extracting subnetworks with good inductive biases that facilitate performance. Supporting this idea, recent work showed that applying IMP to fully connected neural networks (FCNs) leads to the emergence of local receptive fields (RFs), a feature of mammalian visual cortex and convolutional neural networks that facilitates image processing. However, it remains unclear why IMP would uncover localised features in the first place. Inspired by results showing that training on synthetic images with highly non-Gaussian statistics (e.g., sharp edges) is sufficient to drive the emergence of local RFs in FCNs, we hypothesize that IMP iteratively increases the non-Gaussian statistics of FCN representations, creating a feedback loop that enhances localization. Here, we demonstrate first that non-Gaussian input statistics are indeed necessary for IMP to discover localized RFs. We then develop a new method for measuring the effect of individual weights on the statistics of the FCN representations (“cavity method”), which allows us to show that IMP systematically increases the non-Gaussianity of pre-activations, leading to the formation of localised RFs. Our work, which is the first to study the effect of IMP on the statistics of the representations of neural networks, sheds parsimonious light on one way in which IMP can drive the formation of strong inductive biases.
@@ 1 Introduction
Iterative magnitude pruning (IMP) [1] has emerged as a powerful tool for identifying sparse subnetworks (“winning tickets”) that can be trained to perform as well as the dense neural network model from which they are extracted [2, 3]. That IMP, despite its simplicity, is more robust in discovering such winning tickets than more complex pruning schemes [4] suggests that its iterative coarse-graining [5] is especially capable of extracting and maintaining strong inductive biases. This perspective is strengthened by observations that winning tickets discovered by IMP are transferable across related tasks [6, 7, 8, 9, 10, 11, 12, 13] and architectures [14]; that they can outperform dense models on classes with limited data [15]; and that they are less prone to overconfident predictions [16].
@@
The first direct evidence for IMP discovering good inductive biases came from Pellegrini and Biroli [17], who found that the sparse subnetworks extracted by IMP from fully-connected neural networks (FCNs) had local receptive fields (RF) (Fig. 1A). Localised RFs are well-suited for image processing, and are also found in the visual cortex [18, 19] and in convolutional neural networks (CNNs) [20, 21]. Comparing the sparse subnetworks found by IMP and oneshot pruning (Fig. 1B), Pellegrini and Biroli [17] argued that the iterative nature of IMP was essential for refining the local RF structure. However, an understanding of how IMP, a pruning method based purely on the magnitude of the network parameters, is able to yield localised receptive fields remains unknown.
@@
Figure 1: IMP discovers more localized RFs than oneshot magnitude pruning in FCNs. (A) Localized RFs are present after applying IMP for 10 rounds of pruning (each round pruning $s=30\%$ of the remaining weights), leading to a subnetwork with $s=97.2\%$. (B) Noisier, less localized RFs are present in the masks found after oneshot pruning FCNs trained on ImageNet32 to $s=97.2\%$ sparsity. Pruned weights are shown in black and remaining weights are colored by which input channel (red, green, blue) they are connected to. The masks shown correspond to the 120 hidden units with the greatest number of weights remaining [17].
@@
Historically, the study of local RFs has focused on specific features of natural images (e.g., sharp edges), and it has been shown that, with regularization, it is possible for local RFs to emerge in FCNs [22, 23, 24, 25]. Recent work has built upon this, showing that synthetic images with sufficiently strong non-Gaussian statistics can, without any regularization, drive the formation of local RFs in FCNs [26, 27]. Inspired by this, we hypothesize that IMP (Sec. 2.1) is able to discover local RFs (Sec. 2.2) in FCNs by iteratively increasing the non-Gaussian statistics present in its internal representation. This could create a feedback loop, where the amplification of non-Gaussian statistics leads to greater localization, which leads to further increases in non-Gaussian statistics. Because of IMP’s broad success in computer vision, we hypothesize that IMP maximally increases the non-Gaussian statistics, by removing exactly the weights that maximize the “non-Gaussanity” of the internal representations. We formalize our hypothesis below (with kurtosis and pre-activations defined precisely in Sec. 3).
@@
Hypothesis (H): IMP discovers local RFs in FCNs by maximally increasing the kurtosis of the network’s preactivations.
@@
While this is a challenging hypothesis to prove, we provide the following evidence in support of it:
@@
  • Training FCNs on a “Gaussian clone” [28] that matches the original dataset’s first two cumulants (mean and covariance), while containing no higher-order cumulants, we find that IMP does not discover localized RFs. This demonstrates that non-Gaussian statistics are necessary for IMP to discover local RFs in FCNs (Sec. 3.1).
@@
  • Measuring the non-Gaussian statistics present in the preactivations of the FCN, we find that IMP leads to representations with more strongly non-Gaussian statistics, as compared to oneshot pruning (Sec. 3.2). This difference increases with each round of pruning, demonstrating the importance of the iterative nature of IMP.
@@
  • Developing a method that measures the effect individual weights have on the statistics of the FCN representations, we find that IMP removes weights precisely when their pruning would increase the non-Gaussanity of the FCN representations (Sec. 3.3). This suggests that IMP not just increases, but maximizes, the non-Gaussian statistics at each round of pruning.
@@
Collectively, our work provides the first in-depth analysis of how IMP affects the statistics of the internal representations of DNNs, as well as provides insight on how these statistics interact with the iteratively evolving architecture achieved by IMP. Additionally, it highlights the importance of going beyond the assumption of Gaussian features that dominates the current theory of machine learning [29, 30, 31, 32, 33, 34, 35]. Our work motivates further study of how IMP learns parsimonious structure in DNNs, and provides tools (Gaussian data clones and the cavity method) to achieve this goal.
@@ 2.1 Iterative magnitude pruning
Given a neural network $f(\theta;\mathcal{X})$, where $\theta$ are its $N$ parameters (e.g. its weights, biases) and $\mathcal{X}$ is a set of data samples used for training, IMP [1, 2] performs the following iterative pruning procedure. First, $f(\theta,\mathcal{X})$ is trained for $T$ iterations, resulting in $f[\theta(T);\mathcal{X}]$. Then, a mask $m\in\{0,1\}^{N}$ is computed by assigning $m_{i}=0$ if the magnitude of the $i$th weight $\theta_{i}(T)$ is smaller than some fixed threshold $\tau>0$. For network parameters with magnitude greater than $\tau$, $m_{i}=1$. Typically, $\tau$ is set such that $s=1-\frac{1}{N}\sum_{i=1}^{N}m_{i}$, for a desired sparsity $s\in(0,1]$. Having computed the mask, the network parameter values are “rewound” [36, 37] to a previous value, $\theta(t_{\text{rewind}})$, where $t_{\text{rewind}}\ll T$. The original work leveraging IMP to discover winning tickets demonstrated that winning tickets could be found at initialization (i.e. $t_{\text{rewind}}=0$). Subsequent work on has found it necessary to set $t_{\text{rewind}}>0$ [3], particularly when considering architectures larger than LeNet [21] and datasets more complex than MNIST. A common approach is to set $t_{\text{rewind}}\approx 0.01T$. The network $f[\theta(t_{\text{rewind}})\odot m;\mathcal{X}]$ is then trained for $T-t_{\text{rewind}}$ training iterations, where $\odot$ denotes element-wise multiplication.
This train, prune, and rewind procedure is then repeated, with round $n$ of IMP involving training $f[\theta(t_{\text{rewind}})\odot m(n-1);\mathcal{X}]$ and computing the mask $m(n)\in\{0,1\}^{N}$ from the remaining (non-pruned) parameters $\theta(T)\odot m(n-1)$. The masks $m(n)$ are non-trainable parameters, and $\sum_{i=1}^{N}m_{i}(n)<\sum_{i=1}^{N}m_{i}(n-1)$. This is repeated for $N_{\text{IMP}}$ rounds.
@@ 2.2 Local receptive fields
Early work on local receptive fields
Local RFs were first discovered in mammalian primary visual cortex (V1) [18, 19, 49], where individual V1 neurons were found to respond to specific visual features (e.g. the presence of lines) in a local range of visual space. The computational paradigm of breaking input space into local patches, over which a hierarchical representation could be learned, inspired the development of CNNs [50, 21]. Early work studying how local RFs might emerge in artificial neural networks showed that $L_{1}$ regularization of hidden unit activations was capable of driving localization in FCNs [22, 23, 24, 25]. Thus, while sparsity in activations was known to lead to local RFs, it was not until subsequent effort using a variant of LASSO regression ($\beta$-LASSO) on network parameters that it was appreciated that sparsity in the weights could also lead to localized RFs [51]. Unlike this prior literature, our work focuses on the emergence of local RFs without explicit regularization [17, 26, 27].
Quantifying the locality of receptive fields
To quantify the locality of the IMP masks, $m(n)$, we use the following metric. Let $X\in\mathcal{X}$, be a square image with $N_{p}^{2}\in\mathcal{N}$ pixels ($N_{p}$ in the $x-$dimension and $N_{p}$ in the $y-$dimension). Let the locations of two pixels of $X$ be denoted as $z=(x,y)$ and $z^{\prime}=(x^{\prime},y^{\prime})$, where $z,z^{\prime}\in[1,...,N_{p}]\times[1,...,N_{p}]$. Let the relative position of $z$ and $z^{\prime}$ be denoted as $d_{zz^{\prime}}=z-z^{\prime}$. We define the number of pixels in $X$, with relative position $d$, that are connected to unit $i$ in the first hidden layer of an FCN, at IMP round $n$, by
@@
where $\delta(d_{zz^{\prime}}-d)$ is a delta function that is equal to $1$ when $d_{zz^{\prime}}=d$, and 00 otherwise. $S_{i}(d,n)$ acts as a correlation function on the mask associated with unit $i$, $m_{i}(n)$. Pelligrini and Biroli (2022) [17] previously used $S(d,n)$ to visualize the locality of the IMP masks. We further use it to quantify the localization by computing the standard deviations, $\sigma_{x}$ and $\sigma_{y}$, of a two-dimensional Gaussian fit to $S_{i}(d,n)$ (Fig. S1A). These standard deviations act as a measure of correlation length, with the smaller $\sigma_{x}$ and $\sigma_{y}$ are, the more localized the RFs of the IMP masks are. We report $\sigma_{x}$ as the RF width (similar results are found with $\sigma_{y}$). Note that, because the Gaussian fits are applied to this correlation function, $S_{i}(d,n)$, it can be appropriate and capture aspects of local RF width, even in cases where IMP masks do not appear to be Gaussian (Fig. S1). For more details, see Appendix A.
@@ 3.1 Non-Gaussian statistics are necessary for IMP to discover local RFs in FCNs
Key statistical properties of images are encoded in their non-Gaussian statistics.
A key property of natural images is the presence of sharp changes in luminosity, for example at the edges of objects. Edges have long been recognized as a hallmark of natural images [22, 23, 24, 25]. Statistically speaking, these edges manifest themselves at the level of higher-order cumulants (HOCs), which capture the correlations between groups of three or more pixels that cannot be decomposed into products of correlations between smaller groups of pixels. For example, the fourth-order cumulant captures four-pixel correlations that cannot be decomposed into products of pair-wise correlations. HOCs are intrinsically non-Gaussian: in a Gaussian distribution, all HOCs are equal to zero. Hence some of the key features of images, such as the presence of sharp edges, cannot be reproduced using a Gaussian model for images here all cumulants after the second are zero. Seminal work on learning local RFs has emphasized the importance of these non-Gaussian statistics for unsupervised learning [22, 23, 24, 25], while more recent work demonstrated that synthetic images with strongly non-Gaussian statistics can be sufficient to drive local RF formation in FCNs trained on a supervised classification task [26, 27].
@@
Figure 2: Non-Gaussian statistics contain local information in ImageNet32. (A) By maximizing the non-Gaussanity of a lower dimensional representation of the 50,000 validation images from ImageNet32, ICA extracts features, some of which are localized. (B) In contrast, considering only the covariance of the validation images from ImageNet32 leads PCA to extract features that are periodic and thus, non-local.
The most non-Gaussian projections of images are localised
To illustrate why non-Gaussian statistics play a crucial role in the learning of local RFs, we perform independent component analysis (ICA) [52] and principal component analysis (PCA) on images from the ImageNet32 dataset [53], a downsampled version of the original ImageNet dataset (see Appendix B for more details). ICA looks for linear projections, $\lambda=W\cdot X$, which maximize the “non-Gaussianity” of the features $\lambda$. Running this algorithm on ImageNet32 leads to components $W$ that have largely localized features (Fig. 2A). In contrast, running PCA, which only acts on the covariance of the inputs (and thus, only has access to the Gaussian features of the data), yields $W$ with oscillating components (Fig. 2B). Similar results are found when applying ICA and PCA to the full ImageNet dataset. These oscillations provide a spatially distributed and thus, non-local, representation of the images. Therefore, local information is carried primarily by the non-Gaussian statistics of natural images. While both the IMP masks and the ICA components are localized, the exact features they discover differ (Fig. S2). This suggests that IMP is not providing an approximation to ICA. This could be due to the fact that IMP is applied to FCNs trained on an image classification task, as opposed to ICA, which is applied as an image reconstruction task.
@@
IMP does not discover local RFs when applied to FCNs trained on data with only Gaussian statistics. Given the results from Fig. 2, we hypothesize that non-Gaussian statistics are necessary for IMP to find local RFs in FCNs. To test this, we generate a Gaussian clone of ImageNet32 (ImageNet32 sampled from a Gaussian process – “ImageNet32-GP”) [28]. This dataset matches the mean and covariance of ImageNet32, but is explicitly constructed to not contain any higher-order cumulants (Fig. 3A – see Appendix C.1, for details). Unlike when IMP is applied to FCNs trained on ImageNet32, IMP fails to find local RFs in the ImageNet32-GP setting. The masks contain no obvious structure (Fig. 3B) and our metric for localization (Sec. 2.2) shows that the masks found on ImageNet32-GP have a significantly larger correlation length than the masks found on ImageNet32. These results demonstrate that non-Gaussian statistics, in the form of higher-order cumulants, are necessary for local RFs to emerge via IMP.